OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5 A.0 B.1 C.2 D.infinite

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

B.\[\text{Solve}\left[\left\{4 y-2 x=-7,y=5-\frac{x}{2}\right\}\right] \]\[\left\{x\to \frac{27}{4},y\to \frac{13}{8}\right\} \]